# Rethinking Channel Dependence for Multivariate Time Series Forecasting: Learning from Leading Indicators

### 31 Jan 2024 v1

@@ 5.3 Ablation Study
To verify the effectiveness of our designs, we introduce several variants of LightMTS by removing the influence term $\sum\nolimits_{k=1}^{K}{\widetilde{U}}^{(j)}_{k}$ in Eq. (11), removing the difference term $\sum\nolimits_{k=1}^{K}{\widetilde{\Delta}}^{(j)}_{k}$ in Eq. (11), and directly using $V^{(}j)$, $\sum\nolimits_{k=1}^{K}{U}^{(j)}_{k}$ and $\sum\nolimits_{k=1}^{K}{\Delta}^{(j)}_{k}$ in Eq. (11), respectively.
As shown in Figure 4(b), we conduct experiments on these variants with $H$ set to 96, reporting the relative MSE w.r.t. LightMTS on Weather, Electricity, and Traffic datasets. With both the influence and the difference involved, LightMTS considers two kinds of lead-lag relationships and keeps the best performance across the datasets. In contrast, LightMTS w/o influence and LightMTS w/o difference only consider one-sided information of leading indicators, thus showing inferior performance, especially on the Electricity dataset. Furthermore, LightMTS w/o filtering achieves the worst results in all the cases, which fails to adaptively filter out the noise in leading indicators.
@@ 5.4 Hyperparameter Study
Notably, LIFT and LightMTS only introduce two additional hyperparameters, i.e., the number of selected leading indicators $K$ and the number of states $N$. Thus it requires a much little labor to perform hyperparameter selection.
With DLinear as the backbone, we conduct hyperparameter studies on LIFT with $H$ set to 96. As shown in Figure 4(c), LIFT achieves lower MSE with an increasing $K$ on most datasets. Nevertheless, LIFT may well include more noise with a too large $K$ (e.g., on the Wind dataset), resulting in performance degradation. Besides, LIFT cannot enjoy significant improvement with a larger $K$ on the Electricity dataset, where the lead-lag relationships are perhaps more sparse. As for variate states, LIFT achieves lower MSE with an increasing $N$ in most cases. Notably, ignoring the variate states leads to the most significant performance drop on Weather. It is noteworthy that the variates in Weather (e.g., wind speed, humidity, and air temperature) are recorded by diverse sensors, and thus the lead-lag patterns naturally vary with the variate states.
@@ 6 Conclusion
In this work, we rethink the channel dependence in MTS and highlight the locally stationary lead-lag relationship between variates. We propose a novel method called LIFT that efficiently estimates the relationships and dynamically incorporates leading indicators in the frequency domain for MTS forecasting. LIFT can work as a plug-and-play module and is generally applicable to arbitrary forecasting models. We further introduce LightMTS as a lightweight yet strong baseline for MTS forecasting, which keeps similar parameter efficiency with linear models and shows considerable performance compared with state-of-the-art methods. We anticipate that the lead-lag relationship can offer a novel cross-time perspective on the channel dependence in MTS, which is a promising direction for the future development of channel-dependent Transformers or other complex neural networks.
@@ C.1 Dataset Descriptions
Max. Coefficient 0.91 0.88 0.89 0.95 0.92 0.90 0.64 0.56 0.83
We provide the statistics of the nine popular MTS datasets in Table 2. Max. Coefficient represents the average cross-correlation coefficient of the most possible leading indicators on the dataset, where we have dismissed the variates that achieve the largest cross-correlation at a zero leading step. The detailed descriptions are listed as follows.
@@ C.2 Implementation Details
All experiments are conducted on a single Nvidia A100 40GB GPU. The source code will be made publicly available upon publication.
We use the official implementations of all baselines and follow their recommended hyperparameters. Typically, the batch size is set to 32 for most baselines, while PatchTST recommends 128 for Weather. We adopt the Adam optimizer and search the optimal learning rate in {0.5, 0.1, 0.05, 0.01, 0.005, 0.001, 0.0005, 0.0001, 0.00005, 0.00001}. As for LIFT, we continue the grid search with $K$ in {1, 2, 4, 8, 12, 16} and $N$ in {1, 2, 4, 8, 12, 16}. Note that we stop the hyperparameter tuning for consistent performance drop along one dimension of the hyperparameters, i.e., we only conduct the search in a subset of the grid.
@@ Appendix D Additional Experiments
D.1 Paramater Efficiency
We compare the parameters of all baselines on the six datasets. As shown in Figure 6, LightMTS keeps similar parameter efficiency with DLinear, a simple linear model. On average, the parameter size of LightMTS is 1/5 of PatchTST, 1/25 of Crossformer, 1/50 of MTGNN, and 1/70 of Informer and Autoformer. It is noteworthy that a larger $H$ enlarges the gap between PatchTST and LightMTS because PatchTST employs a fully connected layer to decode the $H$-length sequence of high-dimensional hidden states. Although the parameter sizes of Informer and Autoformer are irrelevant to $H$, they are still the most parameter-heavy due to their high-dimensional learning throughout encoding and decoding.
@@ D.1 Paramater Efficiency
D.2 Performance on Other Datasets
@@
D.3 Distribution Shifts
To investigate the dynamic variation of lead-lag relationships, we adopt the proposed lead estimator ($K=2$) to count the lead-lag relationships in the training data and test data. As shown in Figure 7, some leading indicators of a specific variate in the training data cannot keep the relationships in the test data, while the test data also encounters new patterns. GNN-based (Wu et al., 2020) and MLP-based (Li et al., 2023b) methods are susceptible to such distribution shifts of leading indicators due to their static parameter weights in channel-dependent modeling.

### 07 Apr 2024 v4

@@ 5.3 Ablation Study
To verify the effectiveness of our designs, we introduce three variants of LightMTS by removing the influence term $\sum\nolimits_{k=1}^{K}{\widetilde{U}}^{(j)}_{k}$ in Eq. (11), removing the difference term $\sum\nolimits_{k=1}^{K}{\widetilde{\Delta}}^{(j)}_{k}$ in Eq. (11), and directly using $V^{(j)}$, $\sum\nolimits_{k=1}^{K}{U}^{(j)}_{k}$ and $\sum\nolimits_{k=1}^{K}{\Delta}^{(j)}_{k}$ in Eq. (11), respectively.
As shown in Figure 4(b), we conduct experiments on these variants with $H$ set to 96, reporting the relative MSE w.r.t. LightMTS on Weather, Electricity, and Traffic datasets. With both the influence and the difference involved, LightMTS considers two kinds of lead-lag relationships and keeps the best performance across the datasets. In contrast, LightMTS w/o influence and LightMTS w/o difference only consider one-sided information of leading indicators, thus showing inferior performance, especially on the Electricity dataset. Furthermore, LightMTS w/o filter achieves the worst results in all the cases, which fails to adaptively filter out the noise in leading indicators.
@@ 5.4 Hyperparameter Study
Our method introduces merely two additional hyperparameters, i.e., the number of selected leading indicators $K$ and the number of states $N$. Thus it requires a little labor for hyperparameter selection.
With DLinear as the backbone and $H$ set to 96, we study the hyperparameter sensitivity of LIFT. As shown in Figure 4(c), LIFT achieves lower MSE with an increasing $K$ on most datasets. Nevertheless, LIFT may well include more noise with a too large $K$ (e.g., on the Wind dataset), resulting in performance degradation. Besides, LIFT cannot enjoy significant improvement with a larger $K$ on the Electricity dataset, where the lead-lag relationships are perhaps more sparse. As for variate states, LIFT achieves lower MSE with an increasing $N$ in most cases. We observe the most significant performance drop on Weather when ignoring the variate states. It is noteworthy that the variates of Weather (e.g., wind speed, humidity, and air temperature) are recorded by various kinds of sensors, and the lead-lag patterns naturally vary with the variate states.
@@ 6 Conclusion
In this work, we rethink the channel dependence in MTS and highlight the locally stationary lead-lag relationship between variates. We propose a novel method called LIFT that efficiently estimates the relationships and dynamically incorporates leading indicators in the frequency domain for MTS forecasting. LIFT can work as a plug-and-play module and is generally applicable to arbitrary forecasting models. We further introduce LightMTS as a lightweight yet strong baseline for MTS forecasting, which keeps similar parameter efficiency to linear models and shows considerable performance. We anticipate that the lead-lag relationship can offer a novel cross-time perspective on the channel dependence in MTS, which is a promising direction for the future development of channel-dependent Transformers or other complex neural networks.
@@ C.1 Dataset Descriptions
We provide the statistics of the nine popular MTS datasets in Table 2. The detailed descriptions are listed as follows.
@@ C.2 Implementation Details
All experiments are conducted on a single Nvidia A100 40GB GPU. We use the official implementations of all baselines and follow their recommended hyperparameters. Typically, the batch size is set to 32 for most baselines, while PatchTST recommends 128 for Weather. We adopt the Adam optimizer and search the optimal learning rate in {0.5, 0.1, 0.05, 0.01, 0.005, 0.001, 0.0005, 0.0001, 0.00005, 0.00001}. As for LIFT, we continue the grid search with $K$ in {1, 2, 4, 8, 12, 16} and $N$ in {1, 2, 4, 8, 12, 16}. Note that we stop the hyperparameter tuning for consistent performance drop along one dimension of the hyperparameters, i.e., we only conduct the search in a subset of the grid.
@@ Appendix D Additional Experiments
D.1 Efficiency Study
Table 3: Comparison of practical efficiency of PatchTST, Crossformer, and LIFT with $H=720$. MACs are the number of multiply-accumulate operations per sample. The batch size is set to 1.
PatchTST +LIFT Relative additional cost Crossformer +LIFT Relative additional cost
Weather Parameter (M) 4.3 4.7 9.3% 11.7 12.5 7.1%
MACs (G) 0.5 0.5 1.8% 4.4 4.4 0.2%
Memory (MB) 43 46 7.7% 206 209 1.6%
Time (ms) 4.7 6.0 29.6% 37.7 40.6 7.6%
Electricity Parameter (M) 4.3 4.7 9.3% 2.4 2.8 16.7%
MACs (G) 7.1 7.2 1.8% 8.6 8.7 1.4%
Memory (MB) 0.4 0.4 1.2% 873 878 0.6%
Time (ms) 5.1 6.0 17.4% 31.5 33.1 5.1%
Traffic Parameter (M) 4.3 4.7 9.3% 3.2 3.6 12.4%
MACs (G) 19.0 19.4 1.8% 10.6 10.9 3.2%
Memory (MB) 1061 1092 2.9% 1558 1578 1.3%
Time (ms) 5.1 5.8 14.5% 32.7 36.4 11.3%
Solar Parameter (M) 2.0 2.9 41.2% 1.9 2.3 23.4%
MACs (G) 0.9 1.0 5.3% 3.7 3.7 1.6%
Memory (MB) 92 96 4.3% 377 381 1.0%
Time (ms) 4.9 6.0 23.4% 31.9 33.4 5.0%
Wind Parameter (M) 2.0 2.5 20.6% 11.8 12.2 3.4%
MACs (G) 0.2 0.2 5.6% 5.7 5.8 0.2%
Memory (MB) 25 29 13.1% 256 259 1.3%
Time (ms) 4.7 6.1 28.9% 45.2 47.2 4.4%
PeMSD8 Parameter (M) 2.0 2.8 35.8% 2.9 4.0 36.5%
MACs (G) 3.5 3.7 5.6% 13.6 13.8 1.5%
Memory (MB) 319 325 1.9% 1381 1387 0.4%
Time (ms) 5.6 6.0 7.0% 32.4 33.7 4.3%
Following Zeng et al. (2023), we compare the backbones and LIFT by the number of parameters, the number of operations, the GPU memory consumption, and the inference time. As shown in Table 3, LIFT additionally requires an average of 10.7% parameters, 1.7% MACs, 2.5% GPU memory, and 14.2% inference time more than its backbone. It is also noteworthy that our Lead Estimator is non-parametric and can pre-compute estimation only once on the training data, reducing the practical training time.
Moreover, we compare the parameter efficiency of LightMTS and all baselines on the six datasets. As shown in Figure 6, LightMTS keeps similar parameter efficiency with DLinear, a simple linear model. On average, the parameter size of LightMTS is 1/5 of PatchTST, 1/25 of Crossformer, 1/50 of MTGNN, and 1/70 of Informer and Autoformer. It is noteworthy that a larger $H$ enlarges the gap between PatchTST and LightMTS because PatchTST employs a fully connected layer to decode the $H$-length sequence of high-dimensional hidden states. Although the parameter sizes of Informer and Autoformer are irrelevant to $H$, they are still the most parameter-heavy due to their high-dimensional learning throughout encoding and decoding.
@@ D.1 Efficiency Study
D.2 Distribution Shifts
To investigate the dynamic variation of lead-lag relationships, we adopt the proposed lead estimator ($K=2$) to count the lead-lag relationships in the training data and test data. As shown in Figure 7, some leading indicators of a specific variate in the training data cannot keep the relationships in the test data, while the test data also encounters new patterns. GNN-based (Wu et al., 2020) and MLP-based (Li et al., 2023b) methods are susceptible to such distribution shifts of leading indicators due to their static parameter weights in modeling channel dependence.
D.3 Performance on Other Datasets
@@
D.4 Performance of LightMTS
In Table 5, we compare the forecasting errors of LightMTS with all baselines. The lookback length $L$ is set to 336 for LightMTS, PatchTST, DLinear, Crossformer, and MTGNN. We borrow the results of other baselines from the paper of PatchTST. As shown in Table 5, LightMTS achieves comparable performance in all the cases.
Table 5: Performance comparison in terms of forecasting errors. The best results are in bold and the second best are underlined
Models LightMTS PatchTST Dlinear Crossformer MTGNN FEDformer Autoformer Informer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.147 0.198 0.152 0.199 0.176 0.237 0.145 0.209 0.157 0.216 0.238 0.314 0.249 0.329 0.354 0.405
192 0.190 0.238 0.197 0.243 0.220 0.282 0.197 0.264 0.205 0.269 0.275 0.329 0.325 0.370 0.419 0.434
336 0.240 0.278 0.249 0.283 0.265 0.319 0.246 0.309 0.258 0.312 0.339 0.377 0.351 0.391 0.583 0.543
720 0.318 0.332 0.320 0.335 0.323 0.362 0.323 0.364 0.335 0.369 0.389 0.409 0.415 0.426 0.916 0.705
Electricity 96 0.131 0.224 0.130 0.222 0.140 0.237 0.142 0.243 0.139 0.242 0.186 0.302 0.196 0.313 0.304 0.393
192 0.148 0.240 0.148 0.240 0.153 0.249 0.159 0.259 0.160 0.261 0.197 0.311 0.211 0.324 0.327 0.417
336 0.165 0.257 0.167 0.261 0.169 0.267 0.192 0.293 0.193 0.284 0.213 0.328 0.214 0.327 0.333 0.422
720 0.203 0.288 0.202 0.291 0.203 0.301 0.225 0.316 0.242 0.327 0.233 0.344 0.236 0.342 0.351 0.427
Traffic 96 0.386 0.268 0.367 0.251 0.410 0.282 0.519 0.293 0.479 0.289 0.576 0.359 0.597 0.371 0.733 0.410
192 0.405 0.276 0.385 0.259 0.423 0.287 0.522 0.296 0.507 0.307 0.610 0.380 0.607 0.382 0.777 0.435
336 0.417 0.282 0.398 0.265 0.436 0.296 0.530 0.300 0.539 0.314 0.608 0.375 0.623 0.387 0.776 0.434
720 0.444 0.298 0.434 0.287 0.466 0.315 0.584 0.369 0.616 0.352 0.621 0.375 0.639 0.395 0.827 0.466
ETTh1 96 0.311 0.359 0.307 0.358 0.375 0.399 0.423 0.448 0.440 0.450 0.326 0.390 0.510 0.492 0.626 0.560
192 0.341 0.374 0.338 0.374 0.405 0.416 0.471 0.474 0.449 0.433 0.365 0.415 0.514 0.495 0.725 0.619
336 0.369 0.391 0.375 0.399 0.439 0.443 0.570 0.546 0.598 0.554 0.392 0.425 0.510 0.492 1.005 0.741
720 0.407 0.416 0.414 0.421 0.472 0.490 0.653 0.621 0.685 0.620 0.446 0.458 0.527 0.493 1.133 0.845
ETTm1 96 0.285 0.340 0.290 0.342 0.299 0.343 0.315 0.370 0.330 0.388 0.326 0.390 0.510 0.492 0.626 0.560
192 0.325 0.364 0.332 0.369 0.335 0.365 0.348 0.390 0.376 0.419 0.365 0.415 0.514 0.495 0.725 0.619
336 0.365 0.384 0.366 0.392 0.369 0.386 0.414 0.432 0.432 0.461 0.392 0.425 0.510 0.492 1.005 0.741
720 0.422 0.423 0.420 0.424 0.425 0.421 0.511 0.552 0.485 0.488 0.446 0.458 0.527 0.493 1.133 0.845